# On cyclotomic matrices involving Gauss sums over finite fields

### v1

@@ Abstract.
Inspired by the Carlitz’s work on cyclotomic matrices, in this paper, we investigate certain cyclotomic matrices involving Gauss sums over finite fields. For example, let $q=p^{n}$ be an odd prime power with $p$ prime and $n\in\mathbb{Z}^{+}$. Let $\zeta_{p}=e^{2\pi{\bf i}/p}$ and let $\chi$ be a generator of the group of all mutiplicative characters of the finite field $\mathbb{F}_{q}$. For the Gauss sum
@@ 1.2. Backgroud and Motivations
(1.5) $\det C_{p}(\psi)=\begin{cases}(-1)^{(p-1)/(2f)}G_{p}(\psi)^{p-1}/p&\mbox{if}\ 2\nmid f,\\ (-1)^{(p-1)/f}G_{p}(\psi)^{p-1}/p&\mbox{if}\ 2\mid f\ \text{and}\ \chi(-1)=1,\\ (-1)^{(f+2)(p-1)/(2f)}G_{p}(\psi)^{p-1}/p&\mbox{if}\ 2\mid f\ \text{and}\ \chi(-1)=-1,\end{cases}$
@@
For any complex number $x,y$ with Re$(x)>0$ and Re$(y)>0$, it is known that the Gamma function is defined by
@@
by sending the character $\psi$ to the Gauss sum $G_{q}(\psi)$, is indeed a finite field analogue of the Gamma function $\Gamma(\cdot)$. Readers may refer to the survey paper [6] for the detailed introduction on this topic.
@@
If we treat $S_{p}$ as a matrix over $\mathbb{F}_{p}$, then Sun [10, Theorem 1.2] proved that $-\det S_{p}$ is always a nonzero square of $\mathbb{F}_{p}$.
@@
Motivated by the above results, in general, for any positive integer $k\mid q-1$, we define the matrices
@@ 1.3. Main Results
(i) If $n\geq 2$, then $B_{q}(1)$ is singular, i.e., $\det B_{q}(1)=0$. For the case $q=p$ we have
@@
(ii) Suppose that $2\nmid p$. Then $\det B_{q}(2)=0$ whenever $n\geq 2$. Also, for $n=1$ we have
@@ 1.4. Outline of This Paper
We will prove Theorem 1.1 in Section 2. The proofs of Theorems 1.21.3 will be given in Sections 3–4 resepectively. Finally, in Section 5 we will pose some problems for further research.
@@
for any $x\in{\mathcal{O}}_{L}$. One can verify that $\chi_{\mathfrak{p}}$ is a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Also, for any integer $10\leq r<q-$ 1, let
@@
Let notations be as above. Then for any integer $10\leq r<q-$ 1 we have
@@
Let $m$ be a positive integer. We next consider a permutation of $\mathbb{Z}/m\mathbb{Z}$. Let $a\in\mathbb{Z}$ with $\gcd(a,m)=1$. Then the map $x\ {\rm mod}\ m\mapsto ax\ {\rm mod}\ p$ induces a permutation $\tau_{a}(m)$ of $\mathbb{Z}/m\mathbb{Z}$. Lerch [9] determined the sign of this permutation.
@@
Let notations be as above and let $\mathrm{sign}(\tau_{a}(m))$ be the sign of $\tau_{m}(a)$. Then
@@
Hence these ${\bm{v}}_{b}$ are linearly independent over $\mathbb{C}$. Now by Lemma 3.3, these $\lambda_{b}$ (where $b\in\mathbb{F}_{q}^{\times}/U_{k}$) are exactly all the eigenvalues of $C_{q}(k)$.
This completes the proof. ∎
@@
The last equality follows from
$\sum_{b\in\mathbb{F}_{q}^{\times}}{\rm Tr}(b)=0.$
@@
Now we are in a position to prove our last theorem.
Proof of Theorem 1.3. (i) Suppose $n\geq 2$. By Lemma 3.6 there is an element $b\in\mathbb{F}_{q}^{\times}$ such that ${\rm Tr}(b)=0$. For this $b$, by Lemma 4.8, Lemma 3.5 and (3.16), we see that
$\frac{1-q}{q}+\frac{1}{q}\lambda_{b}=\frac{1-q}{q}+\frac{q-1}{q}\zeta_{p}^{{\rm Tr}(b)}=0$
is an eigenvalue of the matrix $[G_{q}(\chi^{-i+j})^{-1}]_{0\leq i,j\leq q-2}$. By Lemma 2.2 we have
$\det B_{q}(1)=\pm\det[G_{q}(\chi^{-i+j})^{-1}]_{0\leq i,j\leq q-2}.$
Hence $\det B_{q}(1)=0$ whenever $n\geq 2$.
For $n=1$, by Lemma 4.8, Lemma 3.5 and (3.16) we have
@@
$=\left(\frac{1-p}{p}\right)^{p-1}\cdot\lim_{x\rightarrow 1}\frac{x^{p}-1}{x-1}$
@@
This proves (i).
(ii) Note that $2\nmid p$. Suppose first that $n\geq 2$. By Lemma 3.6 there is an element $b\in\mathbb{F}_{q}^{\times}$ such that ${\rm Tr}(b)=0$. Similar to the above proof, we see that
$\frac{1-q}{q}+\frac{q-1}{2q}\left(\zeta_{p}^{{\rm Tr}(b)}+\zeta_{p}^{{\rm Tr}(-b)}\right)=0$
is an eigenvalue of the matrix $[G_{q}(\chi^{-2i+2j})^{-1}]_{0\leq i,j\leq(q-3)/2}$. By Lemma 2.2 we have
$\det B_{q}(2)=\pm\det[G_{q}(\chi^{-2i+2j})^{-1}]_{0\leq i,j\leq(q-3)/2}.$
Hence $\det B_{q}(2)=0$ whenever $n\geq 2$.
Now we consider the case that $q=p$ is an odd prime. By Lemma 4.8, Lemma 3.5 and (3.16) again, one can verify that
@@
The last equality follows from
$\prod_{b\in\mathbb{F}_{p}^{\times}}\left(1-\zeta_{p}^{b}\right)=\lim_{x\rightarrow 1}\frac{x^{p}-1}{x-1}=p.$
@@
Acknowledgments We would like to thank Prof. Hao Pan for his steadfast encouragement.

### v3

@@ Abstract.
Inspired by the works of L. Carlitz and Z.-W. Sun on cyclotomic matrices, in this paper, we investigate certain cyclotomic matrices involving Gauss sums over finite fields, which can be viewed as finite field analogues of certain matrices related to the Gamma function.
For example, let $q=p^{n}$ be an odd prime power with $p$ prime and $n\in\mathbb{Z}^{+}$. Let $\zeta_{p}=e^{2\pi{\bf i}/p}$ and let $\chi$ be a generator of the group of all mutiplicative characters of the finite field $\mathbb{F}_{q}$. For the Gauss sum
@@ 1.2. Backgroud and Motivations
(1.5) $\det C_{p}(\psi)=\begin{cases}(-1)^{(p-1)/(2f)}G_{p}(\psi)^{p-1}/p&\mbox{if}\ 2\nmid f,\\ (-1)^{(p-1)/f}G_{p}(\psi)^{p-1}/p&\mbox{if}\ 2\mid f\ \text{and}\ \chi(-1)=1,\\ (-1)^{(f+2)(p-1)/(2f)}G_{p}(\psi)^{p-1}/p&\mbox{if}\ 2\mid f\ \text{and}\ \chi(-1)=-1.\end{cases}$
@@
For any complex number $x$ with Re$(x)>0$, the Gamma function is defined by
@@
by sending the character $\psi$ to the Gauss sum $G_{q}(\psi)$, is indeed a finite field analogue of the Gamma function $\Gamma(\cdot)$. Readers may refer to [6] for the detailed introduction on this topic.
@@
If we treat $S_{p}$ as a matrix over $\mathbb{F}_{p}$, then Sun [11, Theorem 1.2] proved that $-\det S_{p}$ is always a nonzero square of $\mathbb{F}_{p}$.
@@
Motivated by the above results, in general, for any positive integer $k\mid q-1$, we define
@@ 1.3. Main Results
(i) Let $k\mid p-1$ be a positive integer. Then $\det B_{p}(k)\neq 0$. In contrast, if $n\geq 2$, then we have $\det B_{q}(k)=0$.
(ii) For the case $k=1$ we have
@@
(iii) If $p$ is an odd prime, then
@@
Remark 1.2.
For Theorem 1.3(i), the condition $k\mid p-1$ cannot be removed. In fact, with the help of computer we verify that
$\det B_{25}(3)\neq 0.$
Inspired by Theorem 1.3(i) and the above remark, we posed the following conjecture.
Conjecture 1.1.
Let $q=p^{n}$ be a prime power with $p$ prime and $n\geq 2$. Then for any positive integer $11\leq k\leq p-$ 1 with $k\mid q-1$ and $k\nmid p-1$ we have
$\det B_{q}(k)\neq 0.$
Remark 1.3.
We will give some detailed discussions on this conjecture in Section 5.
@@ 1.4. Outline of This Paper
We will prove Theorem 1.1 in Section 2. The proofs of Theorems 1.21.3 will be given in Sections 3–4 respectively. Finally, in Section 5 we will pose some problems for further research.
@@
for any $x\in{\mathcal{O}}_{L}$. One can verify that $\chi_{\mathfrak{p}}$ is a generator of $\widehat{\mathbb{F}_{q}^{\times}}$. Also, for any integer $20\leq r\leq q-$ 2, let
@@
Let notations be as above. Then for any integer $20\leq r\leq q-$ 2 we have
@@
Let $m$ be a positive integer. We next consider a permutation of $\mathbb{Z}/m\mathbb{Z}$. Let $a\in\mathbb{Z}$ with $\gcd(a,m)=1$. Then the map $x\ {\rm mod}\ m\mapsto ax\ {\rm mod}\ m$ induces a permutation $\tau_{a}(m)$ of $\mathbb{Z}/m\mathbb{Z}$. Lerch [9] determined the sign of this permutation.
@@
Let notations be as above and let $\mathrm{sign}(\tau_{a}(m))$ denote the sign of $\tau_{m}(a)$. Then
@@
Hence these ${\bm{v}}_{b}$ are linearly independent over $\mathbb{C}$. Now by Lemma 3.3, these $\lambda_{b}$ (where $b\in\mathbb{F}_{q}^{\times}/U_{k}$) are exactly all the eigenvalues of $C_{q}(k)$. This completes the proof. ∎
@@
The last equality follows from $\sum_{b\in\mathbb{F}_{q}^{\times}}{\rm Tr}(b)=0$.
@@
We also need the following result due to Newman [10, Theorem 1].
Lemma 4.9.
Let $N\geq 2$ be a positive integer and let $\zeta_{N}=e^{2\pi{\bf i}/N}$. Let
$\Omega_{N}=\left\{l\in\mathbb{Z}^{+}:\exists\ c_{1},\cdots,c_{l}\in\mathbb{Z}^{+}\ \text{and}\ 0\leq e_{1}<\cdots<e_{l}\leq n-1\ \text{such that}\ \sum_{j=1}^{l}c_{j}\zeta_{N}^{e_{j}}=0\right\}.$
Then $\min(\Omega_{N})=p(N)$, where $p(N)$ is the least prime divisor of $N$.
Proof of Theorem 1.3. (i) Suppose first $n=1$. The case $k=p-1$ is trivial. Now we consider the case $11\leq k<p-$ 1. Then by Lemma 4.8, every eigenvalue of $[G_{p}(\chi^{-ki+kj})^{-1}]_{0\leq i,j\leq m-1}$ can be written as
$\theta_{b}:=\frac{1-p}{p}+\frac{m}{p}\sum_{y\in U_{k}}\zeta_{p}^{by}=\frac{m}{p}\sum_{y\in U_{k}}\left(\zeta_{p}^{by}-1\right)$
where $b\in\mathbb{F}_{p}^{\times}/U_{k}$. As $k<p-1$, by Lemma 4.9 we have
$\theta_{b}=\frac{m}{p}\left(-k\zeta_{p}^{0}+\sum_{y\in U_{k}}\zeta_{p}^{by}\right)\neq 0.$
This implies that $[G_{p}(\chi^{-ki+kj})^{-1}]_{0\leq i,j\leq m-1}$ is nonsingular. Now by Lemma 2.2 we have
$\det B_{p}(k)=\pm\det\left[\frac{1}{G_{p}(\chi^{-ki+kj})}\right]_{0\leq i,j\leq m-1}\neq 0.$
Suppose now $n\geq 2$. As $k\mid p-1$, we have $U_{k}\subseteq\mathbb{F}_{p}^{\times}$. By Lemma 3.6 there is an element $b\in\mathbb{F}_{q}^{\times}$ with ${\rm Tr}(b)=0$. Similar to the above, we see that
$\theta_{b}$ $=\frac{1-q}{q}+\frac{m}{q}\sum_{y\in U_{k}\subseteq\mathbb{F}_{p}^{\times}}\zeta_{p}^{{\rm Tr}(by)}$
$=\frac{1-q}{q}+\frac{m}{q}\sum_{y\in U_{k}\subseteq\mathbb{F}_{p}^{\times}}\zeta_{p}^{y{\rm Tr}(b)}$
$=0$
is an eigenvalue of $[G_{q}(\chi^{-ki+kj})^{-1}]_{0\leq i,j\leq m-1}$. By Lemma 2.2 again, $B_{q}(k)$ is singular.
(ii) Observe first that
(4.19) $\prod_{b\in\mathbb{F}_{p}^{\times}}\left(1-\zeta_{p}^{b}\right)=\lim_{x\rightarrow 1}\frac{x^{p}-1}{x-1}=p.$
Using this and by Lemma 4.8, Lemma 3.5 and (3.16), we have
@@
(iii) By Lemma 4.8, Lemma 3.5, (3.16) and (4.19) again, and noting that $2\nmid p$, one can verify that
@@
5. Concluding Remarks
In this section we focus on Conjecture 1.1. By Theorem 1.3 and the previous proofs, it is easy to see that $\det B_{q}(k)=0$ if and only if there is an element $b\in\mathbb{F}_{q}^{\times}$ such that
(5.20) $\sum_{y\in U_{k}}\left(\zeta_{p}^{{\rm Tr}(by)}-1\right)=\left(-k\zeta_{p}^{0}+\sum_{y\in U_{k}}\zeta_{p}^{{\rm Tr}(by)}\right)=0.$
As $11\leq k<p-$ 1, by Lemma 4.9 we see that (5.20) holds if and only if ${\rm Tr}(by)=0$ for any $y\in U_{k}$.
Set $U_{k}=\{y_{1},\cdots,y_{k}\}$. By the above discussion, Conjecture 1.1 is equivalent to saying that the system of equations
$\begin{cases}y_{1}x+y_{1}^{p}x^{p}+\cdots+y_{1}^{p^{n-1}}x^{p^{n-1}}=0\\ y_{2}x+y_{2}^{p}x^{p}+\cdots+y_{2}^{p^{n-1}}x^{p^{n-1}}=0\\ \cdots\\ y_{k}x+y_{k}^{p}x^{p}+\cdots+y_{k}^{p^{n-1}}x^{p^{n-1}}=0\end{cases}$
has no nonzero solutions over $\mathbb{F}_{q}$. However, we cannot solve this problem currently.
Acknowledgments We would like to thank Prof. Hao Pan for his steadfast encouragement. This work was supported by the Natural Science Foundation of China (Grant Nos. 12101321 and 12201291).